OpenStudy (anonymous):

which of the following are identities A. cos(x+pi/6)=-sin(x-pi/3) B. 1-tanxtany+sin(x+y)/sinxsiny C. tanx-tany+sin(x-y)/cosxcosy D. tan(x-pi/4)=tanx-1

OpenStudy (anonymous):

cos(x+pi/6)=-sin(x-pi/3) is true since cos(x)=sin(x+pi/2)

OpenStudy (anonymous):

there is two answers

Directrix (directrix):

@marissak Option C is not an equation. Where is the = symbol? C) tanx-tany+sin(x-y)/cosxcosy

OpenStudy (anonymous):

sorry, its suppose to be tanx-tany=sin(x-y)/cosxcosy

Directrix (directrix):

@marissak Same question for option B

OpenStudy (anonymous):

and B is suppose to be 1-tanxtany=sin(x+y)/sinxsiny

Directrix (directrix):

@marissak >> Why do you say this: there is two answers

OpenStudy (anonymous):

@directrix because its a choose all that apply question. they usually have two answers

Directrix (directrix):

I agree with @Prebz on cos(x+pi/6)=-sin(x-pi/3) is true @marissak I don't agree that there must be another answer but there could be.

OpenStudy (anonymous):

Okay thank you

Directrix (directrix):

What I'm wondering is if all the grouping symbols are shown in the options. B) 1-tanxtany=sin(x+y)/sinxsiny Same as: 1 - (tanx)(tany) = [sin(x + y)] / [sin(x)*sin(y)] ?

OpenStudy (anonymous):

i just rechecked them and they are the same as i typed them

Directrix (directrix):

@marissak I am sticking with @Prebz A) cos(x+pi/6)=-sin(x-pi/3) is true FINAL ANSWER (Best final answer I see)

OpenStudy (anonymous):

haha okay thank you so much

Directrix (directrix):

@marissak Let us know the answer

OpenStudy (anonymous):

okay i will. i have 4 more questions to go and then it'll tell me if it was right or wrong